The line width for semiconductor manufacturing is assumed to be normally distributed with a mean of 0.5 micrometer and a standard deviation of 0.05 micrometer. What is the probability that a line width is greater than 0.62 micrometer?


1
Expert's answer
2021-11-15T03:32:07-0500

Mean = 0.5

Standard deviation = 0.05


"P(X>0.62)=1-(Z>\frac{X-\mu}{\sigma})"


"P(X>0.62)=1-(Z>\frac{0.62-0.50}{0.05})=1-(Z>2.4)"


From Z normal distribution table, p-value is:


"P(X>0.62)=1-0.9918=0.0082"


This means there is only 0.82% of the probability that a line width is greater than 0.62 micrometer.